OpenStudy (anonymous):

Show the work to solve |x+3| +7 10

OpenStudy (anonymous):

means \[|x+3|+7=10\]?

OpenStudy (anonymous):

or inequality? if equality subtract 7 to get \[|x+3|=3\] and then solve two equations : \[x+3=3\] and \[x+3=-3\]

OpenStudy (anonymous):

so two answers are \[x=0\] or \[x=-6\]